# Survey of Graph Neural Network for Internet of Things and NextG Networks

### v1

@@ I-A Applications of IoT
Over the past decade, the number of connected IoT devices has grown significantly. There are over 30 billion IoT devices, which are expected to quadruple by 2030 [1]. This can be attributed to its myriad of features such as improved quality of life [2], enhanced decision making [3], safety and security [4], efficiency and automation [5], improved accessibility [6], business innovation and growth [7], environmental sustainability [8]. These features have opened the door to the application of IoT across various domains in everyday life as shown in Figure 1. For example, IoT finds its application in smart home systems [9, 10], smart city [11, 12], smart energy systems [13, 14], smart infrastructure [15, 16, 17], smart logistics [18, 19], smart laboratory [20, 21], smart health [22, 23], smart factory [24, 25], smart campus [26], smart surveillance [27, 28], smart agriculture [29, 30], smart transportation [31, 32].
Figure 2: Vision of GNN-enabled 6G Wireless Network.
@@ I-C Era of Machine Learning
In the recent decade due to the exponential increase in computational resources and availability of digital data, Machine Learning (ML) has been making a significant impact in the research community [44, 45, 46]. This is due to the vast potential of ML techniques such as automation of tasks [47], enhanced user experience [48], resource-efficient big data transfer [49] and intelligent decision making [50]. ML is also expected to enable improved energy efficiency and reduced human intervention [51], secure 6G networks [52], intelligent 6G networks [53]. To this end, researchers have used different machine learning techniques such as Feed forward networks [54], Convolutional Neural Networks (CNN) [55, 56, 57], Gated Recurrent Units (GRU) [58, 59], Echo State Networks (ESN) [60, 61], Long-Short Term memory (LSTM) [62, 63] and even hybrid architectures [64]. Additionally, there are techniques such as Generative Adversarial Networks [65], Autoencoder [66], and a category by itself referred to as Deep Reinforcement Learning [67, 68, 69, 70, 71] that using a machine of these machine techniques as its function approximator. These ML techniques fall under a general category called Neural Networks (NN). NNs also have several neural nodes that are connected and information is propagated between them. NNs have become increasingly popular due to its ability to learn complex patterns and relationships from data. This makes them suitable for a wide variety of tasks such as image and speech recognition, natural language processing, and even playing games.
@@ I-D A Primer on Graph Neural Networks
In addition to the aforementioned learning techniques, there is a relatively new class of neural networks called Graph Neural Networks (GNN) [72, 73]. GNNs are based on graphs that are modeled based on two entities, namely nodes (ie., objects) and edges (ie., relationships between nodes). GNNs are well-suited for capturing complex relationships and dependencies in graph data, which is challenging for traditional neural networks. They can learn representations of nodes and edges based on their local neighborhood information. Second, GNNs can handle graphs of varying sizes and structures, making them applicable to a wide range of real-world problems. They can adapt to irregular and dynamic graphs while regular NNs fail. Third, pre-trained GNN models can be fine-tuned on specific graph-related tasks, enabling transfer learning. In this survey, we aim to provide a detailed study of GNN, discussing its design principles and architecture, state-of-the-art literature that covers how GNN has impacted various domains such as IoT, networking, spectrum awareness, and tactical systems, etc. Through this survey, we hope to make it easier for readers to learn and apply GNN to overcome some of the current research challenges of their respective domains of interest.
GNNs have been widely used to address a variety of network management issues such as resource allocation [74, 75, 76, 77, 78], spectrum sensing and sharing [79, 80], privacy [81], edge computing [82, 83], trajectory optimization [84, 85]. Flex-Net [74] uses GNN to jointly optimize the direction of communication and transmission power with low computational complexity as well as enable scalability and generalization capabilities. A GNN variant called Inductive Graph Neural Network Kriging (IGNNK) is used for secure cooperative sensing [79]. The advantages of IGNNK are that a trained GNN model can be transferred to a new graph structure as well as be used to generate data for virtual sensors for a given location as discussed in [86]. Decentralized management and control is one of the key features of next-generation wireless communication. However, privacy leakage is a major issue. To address this issue, in [81] the authors use GNN to design a privacy-preserving and privacy-guaranteed training algorithm for information exchanges among neighbors in wireless networks. In [82], the authors use a combination of GNN and Reinforcement learning (RL) to optimize the task offloading from IoT devices to mobile nodes that is robust to changes in network topology and mobile node failure. In [85], the authors present a GNN-enabled Advanced Actor-Critic (GNN-A2C) algorithm for aerial-edge IoT (EdgeIoT) system to jointly optimize trajectory and task offloading. With the continuous rise in the use of GNNs for a wide range of applications, it is important to keep track of existing literature works as well as have a clear idea of the challenges and opportunities provided by GNNs. Figure 2 shows our vision of GNN-enabled 6G networks with integrated ORAN architecture.
I-E Scope of the article
@@
TABLE I: Comparison of recent surveys related to GNN.
Figure 3: Overall Organization of the Survey.
  • Give a brief introduction to the Internet of Things and discuss the applications of GNN for IoT Network Intrusion Detection and IoT Data Fusion.
  • Provide an overview of spectrum awareness and elucidate the application of GNN for Radio Frequency Signal Classification and Radio Frequency Spectrum Sensing.
  • Delve into the basics of networking and discuss the applications of GNN for networking focusing on Network and Signal Characteristics Prediction, Routing Optimization, Congestion Control, Mobile Edge Computing, Digital Twin Networks, and Unmanned Aerial Networks.
  • Explore the basics of tactical systems and provide a qualitative discussion of the application of GNN for Tactical Communication and Sensing Systems and Localization
  • Deep-dive into the research opportunities and future research direction of GNN to motivate more exploration in the domains specific to next-generation wireless networks.
@@ E Edges
$\textbf{W}_{i}^{(l)},\textbf{W}_{j}^{(l)}$ Weight Matrices of node $i$, node $j$ at layer $l+1$
@@ D Degree Matrix
$||(\cdot)||^{2}$ Squared Euclidean Distance
@@
TABLE II: List of notations and its definition.
@@
GCNN Graph Capsule Convolution
@@
TABLE III: List of commonly used acronyms and their definitions.
@@
The rest of the survey is organized as follows. We formally introduce GNN by discussing its inception and providing a detailed description of the architecture and types of GNN in Section II. In Section III, we discuss the applications of GNN for the Internet of Things, followed by applications of GNN for spectrum awareness, networking, and tactical systems in Section IV, Section V and Section VI, respectively. Finally, we provide our perspective on the opportunities that could motivate novel research directions utilizing GNN in Section VII before concluding in Section VIII. The overall survey overview is shown in Fig. 3
Figure 4: Examples of different GNN terminologies.
@@ II Graph Neural Networks
In this section, we will provide a detailed introduction to Graph Neural Networks (GNN). To this end, we first present the commonly used terminologies in GNN (and neural networks in general) and their definitions in Section II-A, then we will focus on the architecture of GNN in Section II-B and then discuss different types of GNN variants in Section II-C.
II-A Terminologies and Definition
Before delving into the architecture and types of GNN, it is important to lay a strong foundation for basic terminologies used in Graph Neural Networks. Table II lists all important notations and their definitions. Table III lists all important acronyms and their definitions. Below is the list of important terminologies and their definitions used in GNN and the state-of-the-art literature discussed in later sections (Section III - Section VI).
  • Node. Nodes are the fundamental building blocks of a graph and represent an entity or an element in the graph.
  • Edge. Edges are the connections or relationships between nodes defining how nodes are related to each other.
  • Graph. A graph consists of nodes (blue solid circles in Fig. 4(a)) and edges (solid lines between two blue solid circles in Fig. 4(a)). In general, there are two types of graphs namely (i) Directed graphs and (ii) Undirected graphs as shown in Figs. 4(a) and (b), respectively. In a directed graph, edges have a specific direction, indicating a one-way relationship from one node to another. In an undirected graph, edges have no direction, indicating a two-way relationship where the connection is symmetric.
  • Neighborhood. Neighborhood refers to the set of nodes that are connected to the node of interest as shown in Fig. 4(b).
  • Node Feature. The attribute or properties associated with each node is called the node feature. These features include information about the node’s characteristics, attributes, or the context of the graph. Fig. 4(c) shows an example of node features where 0 and 1 represent the presence or absence of a feature for that node.
  • Edge Feature Similar to the node feature, the attributes or properties associated with each edge is called edge features. The edge feature captures the relationship between two nodes as shown in Fig. 4(d). In this example, the numbers represent the distance between the nodes.
  • Embedding. Embedding refers to the process of representing discrete variables as continuous, low-dimensional vectors. There are 3 types of embedding namely Node Embedding, Edge Embedding, and Graph Embedding.
  • Node Embedding. Node embedding involves learning a vector representation for each node in the graph such that nodes with similar network neighborhoods or connectivity patterns have similar embeddings. These embeddings capture the structural and relational information of nodes in the graph.
  • Edge Embedding. Edge embedding aims to learn vector representations for edges in the graph, capturing the relationship between connected nodes. Edge embeddings can encode information such as edge type, weight, or directionality.
  • Graph Embedding. Graph embedding involves learning a fixed-size vector representation for the entire graph. This representation summarizes the structural and topological properties of the graph, making it useful for tasks such as graph classification, clustering, or graph generation.
  • Aggregation. Aggregation refers to the process of collecting information at the node of interest from all the nodes in the neighborhood. Some of the standard aggregation techniques include Mean Aggregation, Max Aggregation, and Attention-based Aggregation.
  • Propagation. Propagation refers to the process of passing or propagating information within the graph. This process involves iteratively updating the node representations based on the information from their neighbors.
  • Activation Function. Activation Functions allow the defining of a nonlinear family of functions to capture the relationship between the input graph data and its representation. Some examples of non-linear activation functions include Sigmoid or Logistic, Hyperbolic Tangent, Rectified Linear Unit (ReLU), LeakyReLU, Parametric ReLU, Exponential Linear Units, Softmax, Scaled Exponential Linear Unit (SELU) and Gaussian Error Linear Unit (GELU), among others.
    Figure 5: Architecture of GNN depicting the three main stages.
  • Convolution. In traditional CNNs, convolution is referred as the process of sliding a filter (or kernel) over the input data while performing element-wise multiplication between the filter and the input and then summing up the results to produce a feature map. In GNN, convolution is referred as the process of updating the node features based on a transformation of aggregated information from the node’s neighbors.
  • Layer. In GNN (and NNs in general), a layer refers to a single computational step responsible for transforming the node features and propagating information across the graph. There are different types of layers such as input, hidden, readout, and broadcast layers. These layers can have specific functionality and are further divided into convolution layer, message passing layer, and pooling layer, among others.
  • Pooling. Pooling refers to the process of reducing the dimensionality of the graph while retaining important information. While pooling and embeddings have similar functionality, pooling works on a sequence of embeddings. For traditional CNNs, there are different types of pooling such as Max Pooling, Average Pooling, Global Average Pooling, Global Max Pooling, and Fractional Pooling. While for GNN, common pooling techniques include DiffPool [110], HIBPool [111], MinCutPool [112], NMFPool [113], Top-K Pool [114], SAGPool [115].
  • Feature Matrix. The feature matrix represents the 2-D representation of features of all the nodes in the graph. For example, if a graph has $N$ nodes, and each node has $F$ number of features, then the feature matrix F has a dimension of $N\times F$. Examples of node and edge feature matrices are shown in Figs. 4(e) and (f).
  • Adjacency Matrix. The adjacency matrix represents the connection between the nodes in a graph. An example of a Feature Matrix is shown in Fig. 4(g).
  • Degree Matrix. A degree matrix is a diagonal matrix containing information about the degree of each node. Here, the degree of the node indicates the number of terminating edges for a node. An example of a Feature Matrix is shown in Fig. 4(h).
  • Local Information. Local information refers to the information from either an individual node or information aggregated from the neighbors to compute a local graph representation.
  • Global Information. Global information refers to the information gathered by aggregated information from all the nodes in the graph to compute a global representation of the entire graph.
II-B Architecture of Graph Neural Networks

### v2

@@ I-A Applications of IoT
Over the past decade, the number of connected IoT devices has grown significantly. There are over 30 billion IoT devices, which are expected to quadruple by 2030 [1]. This can be attributed to its myriad of features such as improved quality of life [2], enhanced decision making [3], safety and security [4], efficiency and automation [5], improved accessibility [6], business innovation and growth [7], environmental sustainability [8]. These features have opened the door to the application of IoT across various domains in everyday life as shown in Figure 1. For example, IoT finds its application in smart home systems [9, 10], smart city [11, 12], smart energy systems [13, 14], smart infrastructure [15, 16, 17], smart logistics [18, 19], smart laboratory [20, 21], smart health [22, 23], smart factory [24, 25, 26], smart campus [27], smart surveillance [28, 29], smart agriculture [30, 31], smart transportation [32, 33].
@@ I-C Era of Machine Learning
In the recent decade due to the exponential increase in computational resources and availability of digital data, Machine Learning (ML) has been making a significant impact in the research community [45, 46, 47]. This is due to the vast potential of ML techniques such as automation of tasks [48], enhanced user experience [49], resource-efficient big data transfer [50] and intelligent decision making [51]. ML is also expected to enable improved energy efficiency and reduced human intervention [52], secure 6G networks [53], intelligent 6G networks [54]. To this end, researchers have used different machine learning techniques such as Feed forward networks [55], Convolutional Neural Networks (CNN) [56, 57, 58], Gated Recurrent Units (GRU) [59, 60], Echo State Networks (ESN) [61, 62], Long-Short Term memory (LSTM) [63, 64] and even hybrid architectures [65]. Additionally, there are techniques such as Generative Adversarial Networks [66], Autoencoder [67], and a category by itself referred to as Deep Reinforcement Learning [68, 69, 70, 71, 72] that uses a one of these machine techniques as its function approximator. These ML techniques fall under a general category called Neural Networks (NN). NNs also have several neural nodes that are connected and information is propagated between them. NNs have become increasingly popular due to its ability to learn complex patterns and relationships from data. This makes them suitable for a wide variety of tasks such as image and speech recognition, natural language processing, and even playing games.
Figure 2: Vision of GNN-enabled 6G Wireless Network.
@@ I-D A Primer on Graph Neural Networks
In addition to the aforementioned learning techniques, there is a relatively new class of neural networks called Graph Neural Networks (GNN) [73, 74]. GNNs are based on graphs that are modeled based on two entities, namely nodes (ie., objects) and edges (ie., relationships between nodes).
Uniqueness of GNNs. GNNs are well-suited for capturing complex relationships and dependencies in graph data, which is challenging for traditional neural networks. They can learn representations of nodes and edges based on their local neighborhood information. Second, GNNs can handle graphs of varying sizes and structures, making them applicable to a wide range of real-world problems. They can adapt to irregular and dynamic graphs while regular NNs fail. Third, pre-trained GNN models can be fine-tuned on specific graph-related tasks, enabling transfer learning. In this survey, we aim to provide a detailed study of GNN, discussing its design principles and architecture, state-of-the-art literature that covers how GNN has impacted various domains such as IoT, networking, spectrum awareness, and tactical systems, etc. GNNs are the go-to option for complex network control problems and applications with networks that are dynamic and need good generalizability and scalability with growing needs for better connectivity and high quality of service. Through this survey, we hope to make it easier for readers to learn and apply GNN to overcome some of the current research challenges of their respective domains of interest.
TABLE I: Comparison of recent surveys related to GNN.
@@
Advantages and Disadvantages of GNNs. GNNs offer superior flexibility and modeling capabilities for data that is structured as graphs, which RNNs or CNNs are not equipped to handle. GNNs can naturally encode node features, edge attributes, and global graph-level properties through iterative message passing and neighborhood aggregation, allowing the model to learn representations that are aware of both local and global topology. More details about message passing and neighborhood aggregation will be discussed later (Section II-A). In contrast, CNNs operate on fixed grid structures (e.g., images), and RNNs on ordered sequences, making them unsuitable for scenarios where connectivity is irregular, such as wireless mesh networks, sensor networks, or social graphs. Additionally, GNNs are permutation-invariant, that is, the output does not depend on the order of nodes or edges. Furthermore, GNNs can integrate edge features such as link capacity, SNR, latency directly into the computation via attention or gating mechanisms, which is not possible with CNNs or RNNs. From a computational perspective, GNNs are comparatively computationally efficient even for large graphs, unlike CNNs which waste resources on irrelevant zero-padding in sparse spatial inputs. Additionally, advanced GNN variants like Graph Attention Networks (GATs), and Spatio-Temporal GNNs (ST-GCN) enable fine-grained control over information flow across the graph and time, outperforming RNNs in spatio-temporal forecasting tasks (e.g., traffic flow, node failures). More details about different types of GNN will be discussed later in this paper (Section II-B). Finally, GNNs exhibit strong inductive learning capability, meaning a model trained on one graph can generalize to unseen graphs with similar structural patterns which is something CNNs and RNNs struggle with due to their tight coupling with input shape and size.
While GNN demonstrates capability to be applied for a variety of problems, they cannot be arbitrarily used as a solution for tasks that lack an inherent or meaningful graph structure, as forcing data into a graph format without natural relation or dependency can lead to poor performance, overfitting, and unnecessary computational overhead. Let us consider an example of RF spectrum monitoring scenario, where the objective to map the spectrum activity within a geographical area where a number of sensors are deployed. In this scenario, the spectrum activity observed at each sensor can be aggregated to form the global spectrum activity map. In this scenario, the sensors can be considered as nodes, the links between the nodes forms the edges, the sensed spectrum measurements are the node features and the the distance between the nodes are the edge features. However, if we modify this scenario to where there are a number of transmitters deployed within a geographical area and the objective is RF fingerprinting i.e, identifying the transmitter based on the received signal, modeling this problem as a graph with the transmitters as nodes is not ideal since signal from one transmitter (i.e, node) does not provide useful information about another sensor’s fingerprint. The fingerprinting depends only on individual transmitter characteristics, not on any neighboring transmitters. This is a simple example to show how a minor change in the network and the network problem could deem GNN an infeasible option and instead RNNs can be used [98, 99].
General Applications of GNNs. GNNs have been widely used to address a variety of network management issues such as resource allocation [100, 101, 102, 103, 104], spectrum sensing and sharing [105, 106, 107], privacy [108], edge computing [109, 110], trajectory optimization [111, 112]. Flex-Net [100] uses GNN to jointly optimize the direction of communication and transmission power with low computational complexity as well as enable scalability and generalization capabilities. A GNN variant called Inductive Graph Neural Network Kriging (IGNNK) is used for secure cooperative sensing [105]. The advantages of IGNNK are that a trained GNN model can be transferred to a new graph structure as well as be used to generate data for virtual sensors for a given location as discussed in [113]. There is another inductive learning method called GraphSAINT [114] which samples the training graphs into smaller mini-batches. This approach reduces the “neighbor explosion” problem, maintains training efficiency even for deeper networks, and seamlessly supports architectural enhancements like attention mechanisms and skip connections. GraphSAINT has been used for applications such as Botnet Detection [115], Hardware Trojan Detection [116], among others. There is a relatively new variant of GNN called Factor Graph Neural Network (FGNN) [117, 118] that has been used for cooperative combinatorial optimization problems [119], localization [120], route planning [121].
Decentralized management and control is one of the key features of next-generation wireless communication. However, privacy leakage is a major issue. To address this issue, in [108] the authors use GNN to design a privacy-preserving and privacy-guaranteed training algorithm for information exchanges among neighbors in wireless networks. In [109], the authors use a combination of GNN and Reinforcement learning (RL) to optimize the task offloading from IoT devices to mobile nodes that is robust to changes in network topology and mobile node failure. In [112], the authors present a GNN-enabled Advanced Actor-Critic (GNN-A2C) algorithm for aerial-edge IoT (EdgeIoT) system to jointly optimize trajectory and task offloading. One of the important research directions for energy-efficient network management is trajectory optimization, which refers to the set of steps taken by any autonomous system to optimize a given network control problem. These network control problems can be maximizing throughput, minimizing end-to-end delay, IoT data fusion, and swarm UAV deployment. With the continuous rise in the use of GNNs for a wide range of applications, it is important to keep track of existing literature works as well as have a clear idea of the challenges and opportunities provided by GNNs. Figure 2 shows our vision of GNN-enabled 6G networks with integrated ORAN architecture.
TABLE II: List of notations and its definition.
@@ E Edges
$\textbf{W}_{i}^{(l)}$ Trainable weight matrix of node $i$ at layers $l$
$\textbf{W}_{i}^{(l+1)}$ Trainable weight matrix of node $i$ at layers $l+1$
$\textbf{W}_{j}^{(l)}$ Trainable weight matrix of node $j$ at layers $l$
$\textbf{W}_{j}^{(l+1)}$ Trainable weight matrix of node $j$ at layers $l+1$
@@ D Degree Matrix
$(\cdot)||^{2}$ Squared Euclidean Distance
@@
TABLE III: List of commonly used acronyms and their definitions.
@@
GCNN Graph Capsule Convolution Neural Network
@@
I-E Scope of the article
The scope of the article is to survey the applications of GNN for key areas of research in IoT and NextG wireless networks. The applications of GNN for wireless domain is still in its infancy. In this survey we explored and compiled applications of GNN for wireless domain that is worth reviewing to enable maturity of GNN as well as future wireless networks. To this end, we selected IoT, RF Spectrum awareness, networking and tactical systems given these are our area of research focus.Collectively, these form the most important and key problems that needs to be addressed to deliver intelligent, secure, and adaptive connectivity solutions for IoT and nextG networks.
  • In today’s world, all IoT smart applications rely on IoT data fusion to aggregate and process vast amounts of heterogeneous data, facilitating real-time decision-making in domains such as healthcare, smart cities, and industrial automation. As IoT ecosystems expand, ensuring network security becomes crucial to protect the vast amount of data collected from any external attacks, making IoT network intrusion detection essential for identifying cyber threats and safeguarding sensitive data. For this reason in this article we first give a brief introduction to the Internet of Things and discuss the application of GNN for IoT smart applications, IoT data fusion and Intrusion Detection.
  • While small-scale indoor IoT networks may rely on wired network setup, there are many IoT applications that share data to a cloud or other data processing centers wirelessly. This makes it important to efficiently utilize the radio spectrum is critical for sustaining the high-speed, low-latency communication demanded by nextG networks. Radio frequency spectrum sensing and radio frequency signal classification optimize spectrum allocation, minimize interference, and enhance dynamic access to limited spectrum resources, thereby ensuring seamless wireless connectivity. To this end, we provide an overview of spectrum awareness and elucidate the application of GNN for Radio Frequency Spectrum Sensing and Radio Frequency Signal Classification.
  • In addition to spectrum allocation and dynamic access, it is also important to manage the overall network to keep up with growing number of IoT devices. Network and signal characteristics prediction enables proactive adjustments to network parameters, improving routing optimization and congestion control for efficient data transmission. Mobile edge computing enhances real-time processing by decentralizing computation, reducing the load on cloud infrastructure and enabling rapid decision-making for IoT applications. Digital twin networks create virtual replicas of physical systems, which helps in mimicking the real-world application performance and aids in developing and testing virtually before integrating them to the real-world scenario. Additionally, unmanned aerial networks extend connectivity to remote and underserved areas, supporting disaster response and autonomous systems. With this goal in mind, we delve into the basics of networking and discuss the applications of GNN for networking focusing on Network and Signal Characteristics Prediction, Routing Optimization, Congestion Control, Mobile Edge Computing, Digital Twin Networks, and Unmanned Aerial Networks.
  • In mission-critical environments, tactical communication and sensing systems ensure secure and resilient operations, with target recognition and localization playing key roles in defense, surveillance, and autonomous navigation. To further shed some light on this area of research we explore the basics of tactical systems and provide a qualitative discussion of the application of GNN for Tactical Communication and Sensing Systems and Localization.
  • For each section, we provide a table that shows the type of ML (i.e,, GNN models used), the application as well as the dataset or simulator used for performance evaluation. Through these tables we hope to provide a quick start-guide to researchers who are interested in using GNN in their research area of interest.
  • Finally we deep-dive into the research opportunities and future research direction of GNN to motivate more exploration in the domains specific to next-generation wireless networks.
@@ I-E Scope of the article
The rest of the survey is organized as follows. We formally introduce GNN by discussing its inception and providing a detailed description of the architecture and types of GNN in Section II. In Section III, we discuss the applications of GNN for the Internet of Things, followed by applications of GNN for spectrum awareness, networking, and tactical systems in Section IV, Section V and Section VI, respectively. Finally, we provide our perspective on the opportunities that could motivate novel research directions utilizing GNN in Section VII before concluding in Section VIII. List of notations and its definition, and commonly used acronyms are tabulated in Table.II and Table III, respectively. The overall survey overview is shown in Figure 3.
Figure 3: Overall Organization of the Survey.
Figure 4: Architecture of Simple GNN.
Figure 5: Architecture of GNN based on Message Passing Neural Networks.
Figure 6: Examples of GNN graph representation and different matrices.
@@ II Graph Neural Networks
In this section, we will provide a detailed introduction to Graph Neural Networks (GNN). To this end, we first present the architecture of GNN in Section II-A and then discuss different types of GNN variants in Section II-B.
Figure 7: Structure of Graphs with Nodes, Node Features, Edges, and Edge Features. In the example graph structure, the solid blue circle with numbers denote the node and the blue lines in-between denotes the edges.
Figure 8: Representative Architecture of GNN variants.
II-A Architecture of Graph Neural Networks